Circular No. 7162

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 J1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports observations of another apparent Kreutz sungrazing comet that he discovered in images taken with both the C2 and C3 coronagraphs aboard SOHO. There was no tail evident, and on May 8.10 UT the comet’s apparent total V magnitude was 5.8. Astrometric measurements made by Biesecker (and reduced by B. G. Marsden, Smithsonian Astrophysical Observatory), together with orbital elements by Marsden, are given on MPEC 1999-J29.

1999 UT  α_{2000}  δ_{2000}
May 7.804  3^{h}04.^{m}8  +14^{o}56′

SUPERNOVA 1999bx IN NGC 6745 S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999bx was obtained by P. Berlind with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on May 6.5 UT. The low-signal-to-noise spectrum exhibits a blue continuum, with hints of broad Hα, Hβ, and He I emission, indicating that the supernova is probably of type II at an early epoch. Narrow emission lines of Hα, Hβ, and [O III] from the host galaxy yield a recession velocity of 4720 km/s. An earlier spectrum, obtained by E. Howell with the McDonald 2.1-m telescope on Apr. 29 UT, showed only a featureless blue continum.

SUPERNOVA 1999bz IN UGC 8959 Jha et al. also communicate: “Spectra of SN 1999bz were obtained

with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph) by P. Berlind on May 6.3 UT and by M. Calkins on May 8.3. The spectra show a flat continuum with broad (FWHM 12 000 km/s) absorption features, the most prominent of which is observed at 638 nm. Given the NASA/IPAC Extragalactic Database recession velocity for UGC 8959 of 25362 km/s, we identify this feature as Na I. The spectra are, on the whole, similar to a spectrum of the type-Ic SN 1994I at roughly 10 days past maximum light (Filippenko et al. 1995, Ap.J. 450, L11). We thus conclude that SN 1999bz is of type Ic, one to two weeks past maximum light. A spectrum of SN

1999bz is available at http://cfa-www.harvard.edu/cfa/oir/Research/
supernova/RecentSN.html.”

1999 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green